OpenStudy (anonymous):

Please help: Write the equation of a line that passes through the point.. (-1,4) with a slope of -1 in standard form

OpenStudy (anonymous):

y-4=(-1)*(x-(-1)) y-4=-x-1 y=-x+3 standard form: x+y-3=0

OpenStudy (anonymous):

Thank you so much,can you help with my other ones?

OpenStudy (anonymous):

ok